Estimate the mean (show your work) using the data below:

      interval     freq

       0-2         7

       3-5        10

       6-8         3

Show work
Answer #1

Midpoint = lower value + Upper value / 2

For interval 0-2 ,

midpoint = 0 + 2 / 2 = 1

Similarly,

For interval 3 - 5 , midpoint = 4

For interval 6-8 , midpoint = 7

Let x be the midpoint and f be the frequency.

Mean = X * f / f

= (1 * 7 + 4 * 10 + 7 * 3) / (7 + 10 + 3 )

= 3.4